# LASPATED: A Library for the Analysis of Spatio-Temporal Discrete Data (User Manual)

### v2

@@ Abstract
This is the User Manual of the LASPATED library. This library is available on GitHub (at https://github.com/vguigues/LASPATED)) and provides a set of tools to analyze spatio-temporal data. A video tutorial for this library is available on Youtube. It is made of a Python package for time and space discretizations and of two packages (one in Matlab and one in C++) implementing the calibration of the probabilistic models for stochastic spatio-temporal data proposed in the companion paper Guigues et al. (2023).
@@ 4.2 Periodic with elementary time windows of different durations
We can modify the previous time discretization keeping periodic observations but with elementary time windows of different durations. The same method add_time_discretization will be used with the first argument still providing the time unit (’Y’, ’M’, ’W’, ’D’, ’H’, ’m’, or ’S’). The second argument now is a list providing the durations of successive elementary time windows. For instance, in Example LABEL:list:2 which follows, the time unit is ’M’ (month) and the second argument is the list [3, 4, 2, 1, 2] providing successive elementary time windows of 3 months, 4 months, 2 months, 1 month, and 2 months. The third argument is the period, which should be a multiple of the sum of the durations specified in the list of the second argument. In Example LABEL:list:2 the period is 12 months meaning that we count the number of observations for January-February-March in a first group, for April-May-June-July in a second group, for August-September in a third group, for October in a fourth group, and for November-December in a fifth group. In this example, if the last argument was 24 instead of 12 we would have 10 groups of time observations: observations for January-February-March of years 1, 3, 5,$\ldots$, (with the years counted from the first year of historical data) in a first group, for April-May-June-July of years 1, 3, 5,$\ldots$, in a second group, for August-September of years 1, 3, 5,$\ldots$, in a third group, for October of years 1, 3, 5,$\ldots$, in a fourth group, for November-December of years 1, 3, 5,$\ldots$, in a fifth group, for January-February-March of years 2, 4, 6,$\ldots$, in a sixth group, for April-May-June-July of years 2, 4, 6,$\ldots$, in a seventh group, for August-September of years 2, 4, 6,$\ldots$, in a eighth group, for October of years 2, 4, 6,$\ldots$, in a ninth group, and for November-December of years 2, 4, 6,$\ldots$, in a tenth group.
@@ 5.2 Rectangular discretization
Figure 2: Space discretization of a rectangle containing the city of Rio de Janeiro in 100×100=10 00010010010000100$\times$ 100=10\,$0$ 100 = 10 000 rectangles, 4916 of which have nonempty intersection with the city and are represented in the figure.
@@ 5.3 Hexagonal discretization
Figure 3: Space discretization of a rectangle containing the city of Rio de Janeiro in hexagons using Uber library H3 for space discretization in hexagons with scale parameter equal to $7$.
Figure 4: Space discretization of a rectangle containing the city of Rio de Janeiro in hexagons using Uber library H3 for space discretization in hexagons with scale parameter equal to $8$.
@@ 6 Additional discretization features
    If we assume the index set $\mathcal{J}$ for the second discretization $D_{2}$ is partitioned into $k$ index sets $\mathcal{J}_{1}$, $\ldots$,$\mathcal{J}_{k}$: $\mathcal{J}=\sum_{i=1}^{k}\mathcal{J}_{i}$, then LASPATED method add_geo_variable can be invoked with type_geo_variable equal to "area" can also be used to compute $\mathcal{A}_{i\ell}=\displaystyle\sum_{j\in\mathcal{J}_{\ell}}\mathcal{A}(i,j)$ for every $\ell=1,2,\ldots,k$, and $i\in\mathcal{I}$. In Listing LABEL:list:9, we read a shape file soil_use.shp (available on the GitHub page of the project) providing the different land types (we have 16 of them in this data) for the city of Rio de Janeiro. Each elementary cell of this discretization contains one land type only. We then filter the first two land types. It then suffices to pass the corresponding GeoDataFrame object to method add_geo_variable. The object app will contain information about the areas of land types 1 and 2 for every elementary cell of discretization $D_{1}$ (see the next section).
@@ 7 Retrieving and writing discretization data in files
LASPATED spatio-temporal discretizations are stored in the dataframe app.events_data and in the GeoDataFrame app.geo_discretization. Dataframe app.events_data has the same columns as the events dataframe loaded in Section 2, plus time and spatial discretization columns. For each event, the time index discretizations are added with the columns named tdiscr_i, i$=0,$$,T-1$, with T being the number of time index discretizations defined. Custom column names may be provided in the app.add_time_discretization method, with the default column_name parameter. Similarly, the spatial discretization is added with the default name gdiscr. For each event, gdiscr is the index of the region where the event took place. The regions defined in the spatial discretization are fully detailed in the app.geo_discretization dataframe. For each region, the dataframe stores its index, the list of indexes of neighboring regions, the latitude/longitude of the centroid of the region, and columns with its additional discretization features (for instance population, areas of different land types, see Section 6 and Example LABEL:list:9 for these features).
@@ 8.1.1 Calibration of the models for fixed penalties
LASPATED function laspated calibrates the models described in Section 2 of the companion paper Guigues et al. (2023). For this function, fixed penalties $w_{i,j}$ and $W_{G}$ are used and the optimization problems to be solved to calibrate the models are solved with the projected gradient method with Armijo line search along the feasible direction. In this section, we focus on laspated function.
@@
  • nbObservations, which gives the number of observations for every possible combinations of arrival types, time steps, and regions. More precisely, if model=noreg then nbObservations is a 3 dimensional array with nbObservations(c,i,t) the number of observations for arrival type with index c, region with index i, and time window with index t (this is parameter $N_{c,i,t}$ of the model without regressors of Guigues et al. (2023), Section 2.) and if model=reg then nbObservations(c,d,t,i) is the number of observations for arrival type $c$, day or holiday $d$, time window with index t, and region with index i (this is parameter $N_{c,d,t,i}$ of the model with regressors of Guigues et al. (2023), Section 2).
  • nbArrivals: array providing the number of arrivals. More precisely, if model=noreg then this is a 3 dimensional array with nbArrivals(c,i,t) the total number of arrivals observed for time window t, region i, arrival type c (this is parameter $M_{c,i,t}$ of the model without regressors of Guigues et al. (2023), Section 2) and if model=reg then nbArrivals(c,d,t,i) is the number of arrivals for arrival type $c$, day or holiday $d$, time window with index t, and region with index i (this is parameter $M_{c,d,t,i}$ of the model with regressors of Guigues et al. (2023), Section 2);
@@
  • sigma: this is the value of parameter $\sigma$ in the pseudo-code for projected gradient with Armijo line search along the feasible direction given in Section 2 of the companion paper Guigues et al. (2023);
  • iterMax: maximal number of iterations for projected gradient; this is parameter MaxNumberIteration for the pseudo-code in Section 2 of the companion paper Guigues et al. (2023);
  • epsilon: parameter $\varepsilon$ for the models (reg and noreg) in Section 2 of the companion paper Guigues et al. (2023);
@@
  • params.weight: 1-dimensional array with params.weight(i) the weight $W_{G}$ corresponding to time group $G$ with index i in the loss function; see Section 2 of the companion paper Guigues et al. (2023) for the expression of the loss function;
@@
  • params.alpha: this is the common value of weights $w_{ij}$ in the loss function (given in Section 2 of Guigues et al. (2023)) which are equal to params.alpha only when regions with indexes $i$ and $j$ are neighbors;
  • params.delta: parameter $\delta$ in stopping criterion (15) of LASPATED paper Guigues et al. (2023);
@@ 8.1.2 Calibration of penalties
The calibration of parameters $w_{i,j}$ and $W_{G}$ for the loss function of the model without regressors from Section 2 in Guigues et al. (2023)) can be done calling function crossValidation of LASPATED library. This function uses cross validation to estimate a value $\alpha=w_{i,j}=W_{G}$ common to all pairs of neighboring regions $i,j$ and all groups $G$. A call to this function is
@@ 8.1.3 Matlab examples
which builds the data corresponding to the example of Section 6.1 in Guigues et al. (2023) and runs laspated calibration function for the model without regressors for a set of penalizations. It also runs cross validation. The output weight is the best penalization found by cross validation and lambda is the vector of intensities.
@@
which builds the data corresponding to the example of Section 6.2 in Guigues et al. (2023) and runs laspated calibration function for this example.
@@ 8.2.1 Compilation and dependencies
  • Gurobi (https://www.gurobi.com), used only in the model with covariates;
  • xtensor and xtl libraries;
When compiling the code, the user need to pass the location of the header files for laspated, xtl, xtensor and gurobi, as well as the location for the Gurobi libraries and their respective flags. xtl and xtensor requires the C++14 standard and its flag must also be passed. The example below compiles an executable called laspated from a main.cpp that includes laspated. The example considers that the Gurobi solver will be used.
@@
We provide five test models in the source file test_problems.cpp and test_problems.h in the Model_Calibration directory. They were used for assert the quality of the projected gradient algorithm.
@@ 8.2.2 Classes and attributes
  • double EPS: tolerance used in comparisons and in the projection algorithm;
@@
  • double cv_proportion: proportion of samples used in each sub-iteration of the cross validation.
@@
  • distance: distance matrix of zones;
@@
  • which_group: groups of each time index;
@@
  • regressors: regressors array. regressor(j,r) is the value of the j-th regressor for zone r;
@@
                xt::xarray<int> &nb_arrivals, xt::xarray<double> &reg,
@@
  • xt::xarray<double> projection(xt::xarray<double> &x): returns the projection of x into the model feasible region,
  • bool is_feasible(xt::xarray<double> &x): returns true if x is in the feasible set,
@@ 8.2.3 Calibration of the model without regressors
Example LABEL:list:11 calls the calibration method for the model of section 2.1 from the paper. For simplicity, we ommitted the code that initializes all the model attributes. We begin including laspated header and using the laspated namespace. Next, we build a RegularizedModel object m, a Param object param and then run the projected gradient method, storing the result in the xarray lambda.
@@
8 // Load variables nb_observations, nb_arrivals, durations, groups,
9 weights, alphas, distance, type_region, neighbors.
@@ 8.2.4 Cross Validation
Cross validation can also be run for the model without regressors. the cross_validation function has the following signature:
@@
Example LABEL:list:13 shows how to execute the cross validation. The user must set the parameter cv_proportion $\in(0,1]$ in the Param object, build a RegularizedModel object and provide a vector of cv_weights. The cross validation will evaluate how the model performs considering each weight in cv_weights as both the space and time regularization penalties. The return struct contains the time spent performing cross validation, the best weight and the solution of the model considering the best weight. Note that the alphas and weights passed when building m are ignored in the cross validation, being replace with each weight in cv_weights during the cross validation.
@@
8 // After loading variables nb_observations, nb_arrivals, durations, groups,
9 weights, alphas, distance, type_region, neighbors...
@@ 11 Param param;
14 RegularizedModel m(nb_observations, nb_arrivals, durations, groups,
15 weights, alphas, distance, type_region, neighbors, param);
@@ 8.2.5 Calibration of the model with regressors
Example LABEL:list:12 performs the calibration of the model from section 2 of the LASPATED paper Guigues et al. (2023). The same function projected_gradient_armijo_feasible can be used. We pass both the class CovariatesModel and a instance of it to the function. The CovariatesModel objected must be built using the observations, arrivals and covariates (regressors) data. The projected_gradient_armijo_feasible returns the estimated $\beta$ variables found.
@@ 8.2.6 Laspated C++ app
To those unfamiliar with C++, we also provide a laspated binary app that allows running the calibration functions. The user must provide a set of files describing the model and a set of configuration options, either via a configuration file or command line. The options describe the algorithmic parameters and model attributes and file locations. The valid options are the following:
  • double EPS: tolerance used in comparisons and in the projection algorithm;
@@
  • string model_type: the type of model being solved, must be either "reg" (CovariatesModel) or "no_reg" (RegularizedModel);
@@
  • string method: the type of method being used. For the CovariatesModel, only "calibration" is supported (running the projected gradient algorithm once). For RegularizedModel, can be either "calibration" or "cross_validation";
@@
  • string algorithm: the algorithm used. Can be either "feasible" for projected gradient along the feasible region or "boundary" for the projected gradient along the boundary, however "cross_validation" only supports the "feasible" algorithm;
@@
  • string info_file: path of file containing general information about the model. See below for a precise description of the format.
@@
  • arrivals_file: path of file containing the samples of arrivals file.
@@
  • neighbors_file: path of file containing description of each zone, including its type, its covariates, and neigborhoods.
@@
  • alpha_regions_file: path of file containing the space regularizing $R\times R$ matrix.
@@
  • time_groups_file: path of file containing a description of period groups and their temporal regularization penalties.
  • double duration: The duration of each period.
@@
  • cv_weights_file: path of file containing the weights to be used in cross validation.
@@
Note that to sue the CovariatesModel, the user only needs to specify the info, arrivals, and neighbors files. Additionally, for the RegularizedModel, the user must specify the alpha_regions_file and time_groups_file. When performing cross validation, the cv_weights_file is also needed.
@@
The info_file must contain two lines. The first line must contain, separated by spaces, the number of periods $T$ per day, the number of week days $G$, the number of zones $R$, the number of arrival types $C$, the number of covariates $J$ and the number of holidays $H$ on each year. The second line contains, for each week day and holiday, the number of observations $N_{g}$ for that day considered on the samples. Values of index $g$ from 0 to 6 are defined as the weekdays while the holidays are defined as $\{7,8,...,7+H\}$.
@@
The arrivals_file contains the samples observed, where each line describes a sample. Each line contains, separated by spaces, the period $t\in[0,T)$, the day $g\in[0,G)$, the zone $r\in[0,R)$, the arrival type $c\in[0,C)$, the sample index $j\in[0,N_{g})$, the number of arrivals observed, and a flag indicating whether the day is a holiday or not.
@@
The neighbors_file contains the description of each zone. Each line contains the zone index, two values representing coordinates of the zone centroid, a number indicating the zone’s type, J values corresponding to each covariate, followed by a set of pairs of values indicating the index and the distance of each neighbor of the current zone.
@@
The alpha_regions_file describe the space regularizations terms. It contains a $R\times R$ matrix separated by spaces, where entry $(i,j)$ corresponds to the weight included in the objective function of the RegularizedModel if zones $i$ and $j$ are neighbors.
@@
The time_groups_file contains description of groups of periods and their corresponding penalties. The first line must have the number $\mathcal{G}$ of time groups. Then, the subsequent $D*T$ lines must have the group index of each period. Then the subsequent $\mathcal{G}$ lines may contain the weight of each group; if the method is set to "cross_validation", the reading of time weights is skipped.
@@
The cv_weights_file contains the weights to be used in cross validation, separated by spaces. Each weight must non-negative.
@@
The LASPATED app is only supported in Linux operating systems. The user must have the GCC compiler and the Boost library. Libraries xtl and xtensor are also used and provided with the In order to use the CovariatesModel, the user must have Gurobi installed. The directory Model_Calibration/Cpp contains a Makefile that is used to compile the app with:
@@
The Makefile uses GUROBI_HOME to find the Gurobi directory, GUROBI_LIB_VER to find the Gurobi C library version and INCLUDE_BOOST to find the Boost directory. The values displayed above are actually the default values set in the Makefile, so if Boost and Gurobi are installed in those locations, then just the make command will compile the code. Once the laspated app is successfully comnpiled, it can be run with:
@@
    \$ ./laspated -f config --option_1=val_1 --option_2=val_1 ...
@@
After running the laspated app, the output is saved in the path provided by option output_file. If the RegularizedModel was used, the output contains on each line the indexes $c\in[0,C)$, $r\in[0,R)$, $t\in[0,D*T)$, and the intensity $\lambda_{crt}$ estimated by LASPATED. If CovariatesModel is used, then each line correspond to indexes $c\in[0,C)$, $d\in[0,D)$, $t\in[0,T)$, $j\in[0,J)$, and the parameter $\beta_{cdtj}$ estimated by LASPATED.

### v3

@@ Abstract
This is the User Manual of the LASPATED library. This library is available on GitHub (at https://github.com/vguigues/LASPATED)) and provides a set of tools to analyze spatio-temporal data. A video tutorial for this library is available on Youtube. It is made of a Python package for time and space discretizations and of two packages (one in Matlab and one in C++) implementing the calibration of the probabilistic models for stochastic spatio-temporal data proposed in the companion paper laspatedpaper.
@@ 4.2 Periodic with elementary time windows of different durations
[…]
@@ 5.2 Rectangular discretization
Figure 2: Space discretization of a rectangle containing the city of Rio de Janeiro in 100×100=10 $00\times 10$0=10\,000 rectangles, 4916 of which have nonempty intersection with the city and are represented in the figure.
@@ 5.3 Hexagonal discretization
Figure 3: Space discretization of a rectangle containing the city of Rio de Janeiro in hexagons using Uber library H3 for space discretization in hexagons with scale parameter equal to 77.
Figure 4: Space discretization of a rectangle containing the city of Rio de Janeiro in hexagons using Uber library H3 for space discretization in hexagons with scale parameter equal to 88.
@@ 6 Additional discretization features
    If we assume the index set $\mathcal{J}$ for the second discretization $D_{2}$ is partitioned into $k$ index sets $\mathcal{J}_{1}$, $\ldots$$\mathcal{J}_{k}$: $\mathcal{J}=\sum_{i=1}^{k}\mathcal{J}_{i}$, then LASPATED method add_geo_variable can be invoked with type_geo_variable equal to "area" can also be used to compute $\mathcal{A}_{i\ell}=\displaystyle\sum_{j\in\mathcal{J}_{\ell}}\mathcal{A}(i,j)$ for every $\ell=1,2,\ldots,k$, and $i\in\mathcal{I}$. In Listing LABEL:list:9, we read a shape file soil_use.shp (available on the GitHub page of the project) providing the different land types (we have 16 of them in this data) for the city of Rio de Janeiro. Each elementary cell of this discretization contains one land type only. We then filter the first two land types. It then suffices to pass the corresponding GeoDataFrame object to method add_geo_variable. The object app will contain information about the areas of land types 1 and 2 for every elementary cell of discretization $D_{1}$ (see the next section).
@@ 7 Retrieving and writing discretization data in files
LASPATED spatio-temporal discretizations are stored in the dataframe app.events_data and in the GeoDataFrame app.geo_discretization. Dataframe app.events_data has the same columns as the events dataframe loaded in Section 2, plus time and spatial discretization columns. For each event, the time index discretizations are added with the columns named tdiscr_i, i=0,=0,,T1,T-1, with T being the number of time index discretizations defined. Custom column names may be provided in the app.add_time_discretization method, with the default column_name parameter. Similarly, the spatial discretization is added with the default name gdiscr. For each event, gdiscr is the index of the region where the event took place. The regions defined in the spatial discretization are fully detailed in the app.geo_discretization dataframe. For each region, the dataframe stores its index, the list of indexes of neighboring regions, the latitude/longitude of the centroid of the region, and columns with its additional discretization features (for instance population, areas of different land types, see Section 6 and Example LABEL:list:9 for these features).
@@ 8.1.1 Calibration of the models for fixed penalties
LASPATED function laspated calibrates the models described in Section 2 of the companion paper laspatedpaper. For this function, fixed penalties $w_{i,j}$ and $W_{G}$ are used and the optimization problems to be solved to calibrate the models are solved with the projected gradient method with Armijo line search along the feasible direction. In this section, we focus on laspated function.
@@
  • nbObservations, which gives the number of observations for every possible combinations of arrival types, time steps, and regions. More precisely, if model=noreg then nbObservations is a 3 dimensional array with nbObservations(c,i,t) the number of observations for arrival type with index c, region with index i, and time window with index t (this is parameter $N_{c,i,t}$ of the model without regressors of laspatedpaper, Section 2.) and if model=reg then nbObservations(c,d,t,i) is the number of observations for arrival type $c$, day or holiday $d$, time window with index t, and region with index i (this is parameter $N_{c,d,t,i}$ of the model with regressors of laspatedpaper, Section 2).
  • nbArrivals: array providing the number of arrivals. More precisely, if model=noreg then this is a 3 dimensional array with nbArrivals(c,i,t) the total number of arrivals observed for time window t, region i, arrival type c (this is parameter $M_{c,i,t}$ of the model without regressors of laspatedpaper, Section 2) and if model=reg then nbArrivals(c,d,t,i) is the number of arrivals for arrival type $c$, day or holiday $d$, time window with index t, and region with index i (this is parameter $M_{c,d,t,i}$ of the model with regressors of laspatedpaper, Section 2);
@@
  • sigma: this is the value of parameter $\sigma$ in the pseudo-code for projected gradient with Armijo line search along the feasible direction given in Section 2 of the companion paper laspatedpaper;
  • iterMax: maximal number of iterations for projected gradient; this is parameter MaxNumberIteration for the pseudo-code in Section 2 of the companion paper laspatedpaper;
  • epsilon: parameter $\varepsilon$ for the models (reg and noreg) in Section 2 of the companion paper laspatedpaper;
@@
  • params.weight: 1-dimensional array with params.weight(i) the weight $W_{G}$ corresponding to time group $G$ with index i in the loss function; see Section 2 of the companion paper laspatedpaper for the expression of the loss function;
@@
  • params.alpha: this is the common value of weights $w_{ij}$ in the loss function (given in Section 2 of laspatedpaper) which are equal to params.alpha only when regions with indexes $i$ and $j$ are neighbors;
  • params.delta: parameter $\delta$ in stopping criterion (15) of LASPATED paper laspatedpaper;
@@ 8.1.2 Calibration of penalties
The calibration of parameters $w_{i,j}$ and $W_{G}$ for the loss function of the model without regressors from Section 2 in laspatedpaper) can be done calling function crossValidation of LASPATED library. This function uses cross validation to estimate a value $\alpha=w_{i,j}=W_{G}$ common to all pairs of neighboring regions $i,j$ and all groups $G$. A call to this function is
@@ 8.1.3 Matlab examples
which builds the data corresponding to the example of Section 6.1 in laspatedpaper and runs laspated calibration function for the model without regressors for a set of penalizations. It also runs cross validation. The output weight is the best penalization found by cross validation and lambda is the vector of intensities.
@@
which builds the data corresponding to the example of Section 6.2 in laspatedpaper and runs laspated calibration function for this example.
@@ 8.2.1 Compilation and dependencies
  • Gurobi (https://www.gurobi.com), an optimization tooblox used only in the model with covariates;
  • xtensor and xtl libraries.
When compiling the code, the user needs to pass the location of the header files for laspated, xtl, xtensor and gurobi, as well as the location for the Gurobi libraries and their respective flags. xtl and xtensor require the C++14 standard and its flag must also be passed. The example below compiles an executable called laspated from a main.cpp that includes laspated. The example considers that the Gurobi solver will be used.
@@
We provide five test models in the source files test_problems.cpp and test_problems.h in the Model_Calibration directory. They were used to assert the quality of the projected gradient algorithm.
@@ 8.2.2 Classes and attributes
  • double EPS: tolerance used in the projection algorithm;
@@
  • relax_empirical_fix: if set to true, relaxes the constraint that makes the sum of estimated rates per class equal to the empirical rate by class.
@@
  • distance: distance matrix (distance between zones);
@@
  • which_group: groups of each time index. which_group[d][t] is the index of the time group corresponding to day d and time interval t within day d;
@@
  • durations: duration, in hours, for each time index $(d,t)$.
  • regressors: array of regressor values. regressor(j,r) is the value of the j-th regressor for zone r;
@@
                xt::xarray<int> &nb_arrivals,
                xt::xarray<double>& a_durations, xt::xarray<double> &reg,
@@
  • xt::xarray<double> projection(xt::xarray<double> &x): returns the projection of x onto the set defined by the constraints of the model with covariates,
  • bool is_feasible(xt::xarray<double> &x): returns true if x is in the feasible set defined by the constraints of the model with covariates and false otherwise,
@@ 8.2.3 Calibration of the model without regressors
Example LABEL:list:11 calls the calibration method for a model without covariates. For simplicity, we omitted the code that initializes all the model attributes. We begin including laspated header and using the laspated namespace. Next, we build a RegularizedModel object m, a Param object param and then run the projected gradient method, storing the result in the xarray lambda.
@@
8 // Load variables nb_observations, nb_arrivals, durations,
9 // groups, weights, alphas, distance, type_region, neighbors.
@@ 8.2.4 Cross Validation
Cross validation can also be run for the model without regressors. This function allows us to choose the penalization parameters of the model without regressors. The corresponding cross_validation function has the following signature:
@@
Example LABEL:list:13 shows how to execute the cross validation. The user must set the parameter cv_proportion $\in(0,1]$ in the Param object, build a RegularizedModel object and provide a vector of cv_weights. The cross validation will evaluate how the model performs considering each weight in cv_weights as both the space and time regularization penalties. The returned struct contains the time spent performing cross validation, the best weight and the solution of the model considering the best weight. Note that the alphas and weights passed when building m are ignored in the cross validation, being replaced with each weight in cv_weights.
@@
8 // After loading variables nb_observations, nb_arrivals,
9 // durations, groups,weights, alphas, distance,
10 // type_region, neighbors
@@ 8.2.5 Calibration of the model with regressors
Example LABEL:list:12 performs the calibration of a model with covariates. Function projected_gradient_armijo_feasible can still be used for projections. We pass both the class CovariatesModel and an instance of this class to the function. The CovariatesModel object must be built using the observations, arrivals and covariates (regressors) data. The projected_gradient_armijo_feasible returns the estimated $\beta$ variables.
@@ 8.2.6 Laspated C++ app
To those unfamiliar with C++, we also provide a laspated binary app that allows running the calibration functions. The user must provide a set of files describing the model and a set of configuration options, either via a configuration file or command line. The options describe the algorithmic parameters, model attributes, and file locations. The valid options are the following:
  • double EPS: tolerance used in the projection algorithm;
@@
  • double cv_proportion: proportion of samples used in each iteration of cross validation;
@@
  • string model_type: the type of model being solved, must be either "reg" (for the model with covariates) or "no_reg" (for the model without covariates);
@@
  • string method: the type of method being used. For the model with covariates, only "calibration" is supported (running the projected gradient algorithm once). For the model without covariates, can be either "calibration" or "cross_validation";
@@
  • string algorithm: the algorithm used. Can be either "feasible" for Armijo search along the feasible direction or "boundary" for Armijo search along the boundary. However "cross_validation" only supports the "feasible" option;
@@
  • string info_file: path of file containing general information about the model. See below for a precise description of the corresponding file.
@@
  • arrivals_file: path of file containing the samples of arrivals. See below for a precise description of the corresponding file.
@@
  • neighbors_file: path of file containing the description of each zone, including its type, its covariates, and its neighbors. See below for a precise description of the corresponding file.
  • durations_file: path of file containing the durations for each time index. See below for a precise description of the corresponding file.
@@
  • alpha_regions_file: path of file containing the space penalizations $R\times R$ matrix. See below for a precise description of the corresponding file.
@@
  • time_groups_file: path of a file containing a description of time groups and the corresponding penalties in the regularized model. See below for a precise description of the corresponding file.
  • double duration: array with the duration of each period.
@@
  • cv_weights_file: path of a file containing the weights to be used in cross validation. See below for a precise description of the corresponding file.
  • relax_empirical_fix: flag that, when set, relaxes the constraint that makes the sum of estimated rates per class equal to the empirical rate by class.
@@
Note that to use the model with covariates, the user only needs to specify the info, arrivals, neighbors, and duration files. Additionally, for the model without regressors, the user must specify the files alpha_regions_file and time_groups_file. When performing cross validation, the file cv_weights_file is also needed.
@@
The file info_file must contain two lines. The first line must contain, separated by spaces, the number of periods $T$ per day, the number of week days $G$, the number of zones $R$, the number of arrival types $C$, the number of covariates $J$ and the number of holidays $H$ for each year. The second line contains, for each day of the week and holiday, the number of observations $N_{g}$ for that day or holiday. Values of index $g$ from 0 to 6 are used for days of the week while holidays have indexes in $\{7,8,...,7+H\}$.
@@
The file arrivals_file contains the samples observed, where each line describes a sample. Each line contains, separated by spaces, the period $t\in[0,T)$, the day $g\in[0,G)$, the zone $r\in[0,R)$, the arrival type $c\in[0,C)$, the sample index $j\in[0,N_{g})$, the number of arrivals observed, and a flag indicating whether the day is a holiday or not.
@@
The file neighbors_file contains the description of each zone. Each line contains the zone index, two values representing the coordinates of the zone centroid, a number indicating the zone’s land type, J values corresponding to the values of all $J$ covariates, followed by a set of pairs of values (index,distance) indicating the distances of the considered zone to its neighbors (identified by their indexes).
@@
The file alpha_regions_file describes the space penalization parameters. It contains a $R\times R$ matrix (with entries separated by spaces) where entry $(i,j)$ corresponds to the weight in the objective function of the model without covariates if zones $i$ and $j$ are neighbors.
@@
The file time_groups_file contains the description of time groups and their corresponding penalties. The first line must have the number $\mathcal{G}$ of time groups. The subsequent DTD*T lines must have the group index of each period. Then the subsequent $\mathcal{G}$ lines may contain the weight of each group. If the method is set to "cross_validation", the reading of time weights is skipped.
@@
File cv_weights_file contains the weights to be used in cross validation, separated by spaces. Each weight must be non-negative.
The durations file contains the duration of each time index, in hours. The first line contains, separated by spaces, the number of day indexes and the number of time indexes.
@@
The LASPATED app is only supported on Linux operating systems. The user must have the GCC compiler and the Boost library. Libraries xtl and xtensor are also used and provided. In order to use the model with covariates, the user must have Gurobi installed. The directory Model_Calibration/Cpp contains a Makefile that is used to compile the app with:
@@
The Makefile uses GUROBI_HOME to find Gurobi directory, GUROBI_LIB_VER to find the Gurobi C library version and INCLUDE_BOOST to find the Boost directory. The values displayed above are actually the default values set in the Makefile, so if Boost and Gurobi are installed in those locations, then the make command will compile the code. Once laspated app is successfully compiled, it can be run with:
@@
    \$ ./laspated -f --model_type=no_reg --method=calibration
        --info_file=info.txt --arrivals_file=arrivals.txt
        --neighbors_file=neighbors.txt
        --alpha_regions_file=alpha_regions.txt
        --time_groups_file=time_groups.txt
        --durations_file=durations.txt
        --output_file=output.txt
@@
After running laspated app, the output is saved in the path provided by option output_file. If the model without covariates is used, the output contains on each line the indexes $c\in[0,C)$, $r\in[0,R)$, $t\in[0,D*T)$, and the intensity $\lambda_{crt}$ estimated by LASPATED. If the model with covariates is used, then each line correspond to indexes $c\in[0,C)$, $d\in[0,D)$, $t\in[0,T)$, $j\in[0,J)$, and the parameter $\beta_{cdtj}$ estimated by LASPATED.